8(t-3) + 4t= 6(2t+1)-10
Answer by funmath(2933)

8(t-3) + 4t= 6(2t+1)-10 First simplify both sides following the order of operation. (PEMDAS, let me know if you don't know what that is.)
8t+8(-3)+4t=6(2t)+6(1)-10
8t-24+4t=12t+6-10
12t-24=12t-4 Now try to get your t's on one side of the = sign, within the rules of algebra.
-12t+12t-24=-12t+12t-4
-24=-4 Your variables were eliminated and -24 does not = -4, therefore there is no solution.
Happy Calculating!!!